OpenStudy (anonymous):

how would i simplify the difference (2w^2-8w-7)-(7w^2+6w-4) ?

OpenStudy (anonymous):

how many w^2 you got?

OpenStudy (anonymous):

i see 2-7 w^2s how many w we you got?

OpenStudy (anonymous):

ther are two w^2s

OpenStudy (anonymous):

there are 2 w^2 in the left term, i see 7 w^2 in the right term, but we are subtracting the righ tterm 2w^2 - 7w^2 = (2-7) w^2 in total

OpenStudy (anonymous):

there are -8 w in the left term, i see 6 w in the right term ... but we are subtracting the right term -8w -6w = (-8-6)w in total all thats left is the numbers

OpenStudy (anonymous):

i see a -7 on the left, a -4 on the right ... -7--4 = ?

OpenStudy (anonymous):

this is the problem (2w^2-8w-7)-(7w^2+6w-4)

OpenStudy (anonymous):

yes, that is what you posted to start with and what i have been addressing .... now we can make this a complete circle if you simply start at the first post.

OpenStudy (anonymous):

okay do you or do you not know the answer to this problem ?

OpenStudy (anonymous):

i know the answer, you need the practice. this isnt a free answering service we are running here, it is a study site to help people learn how to work the material.

OpenStudy (anonymous):

@mathmale any ideas?

OpenStudy (mathmale):

@janeria : Your question, "do you or do you not know the answer to this problem?" is uncalled for, unfair and in violation of the OpenStudy Code of Conduct. guest.100 was trying to guide you through the solution of this problem (not to give you the answer). Please be more considerate in the future. Read the Code of Conduct to see what it has to say about giving and receiving "answers."

OpenStudy (mathmale):

(2w^2-8w-7)-(7w^2+6w-4) could be rewritten in clearer format as\[(2w^2-8w-7)-(7w^2+6w-4)\] I'd suggest that you rewrite this again in column form and then combine the quantities in each column: 2w^2-8w-7 -(7w^2+6w-4) ------------ I now ask you, @janeria, to re-write this by distributing that (-) sign over everything inside parentheses. Write your result here, please.

OpenStudy (mathmale):

@janeria : The ball's now in your court. Someone will respond if y ou will now do your part. Please see the guidelines I gave you, above.